A wave is travelling in the negative direction has a frequency of 40 Hz ,a wavelength of 3cm and an amplitude of 0.6cm.write the wave function y(x,t) given that y=0.6cm at x =0 and t =0.


1
Expert's answer
2021-03-01T12:40:14-0500

The general equation of the wave function can be written as follows:


"y(x,t)=Asin(\omega t \mp kx),"

here, "A" represents the amplitude of the wave, "\omega" represents the angular frequency of the wave, "k"represents the wave number of the wave, sign minus corresponds to the wave traveling in +"x"-direction and sign plus correspond to the wave traveling in -"x"-direction.

Let's find the angular frequency and wave number of the wave:


"\omega=2\pi f=2\pi\cdot40\ Hz=251.2\ \dfrac{rad}{s},""k=\dfrac{2\pi}{\lambda}=\dfrac{2\pi}{3\cdot10^{-2}\ m}=209\ m^{-1}."

Finally, we can write the wave function y(x,t):


"y(x,t)=(0.006\ m)sin(251.2t+209x)."